OpenStudy (anonymous):

A boat heads upstream a distance of 30 miles in the mission river whose current is running at 5 miles per hour. If the trip back takes an hour less, what was the speed of the boat still in the water

OpenStudy (anonymous):

DistanceAB = DistanceBA = 30 miles rateRIV - rate of the river = 5 miles per hour rateBOAT - rate of the boat in still water timeAB = time traveled by boat upstream timeBA = time traveled by boat downstream (1) timeBA = timeAB - 1 (2) rateBOAT - 5 = 30/timeAB (3) rateBOAT + 5 = 30/timeBA Substitute (1) in (3) for timeBA, and solve (2) and (3) simultaneously... rateBOAT^2 - 325 = 0 rateBOAT = square root (325) = 18.03 miles per hour is the answer.... To check, use (2) and (3).... timeAB = 2.3 hours timeBA = 1.3 hours, less 1 hr than 2.3 hours.... :)

OpenStudy (gorv):

let speed of boat =x speed of boat in upstream=x-5 (boat is going against the current so it will reduce boat speed) distance=30 miles time upstream=30/(x-5) in downstream that is during back...current will speed up the boat so speed=x+5 distance=30 miles time in coming back=30/(x+5)=time upstream -1hour \[30/(x+5)=30/(x-5) -1\]

OpenStudy (gorv):

30/(x+5)=(30-1*(x-5))/(x-5) 30/(x+5)=(30-x+5)/(x-5) 30*(x-5)=(35-x)*(x+5) 30x-150=35x+175-x^2-5x 30x-175=30x+175-x^2 x^2=300 x=300^(1/2)=17.32miles/hour

OpenStudy (gorv):

sorry 30x-150=35x+175-x^2-5x 30x=30x+175+150-x^2 x^2=325 x=18.02miles/hour

OpenStudy (anonymous):

... thanks for proving my solution... :)